OpenStudy (dls):

A simple harmonic oscillator has an amplitude A and Time period T.The time period required by it to travel from x=a to x=a/2 is? A)T/6 B)T/4 C)T/3 D)T/2

OpenStudy (dls):

 maybe smt like this

OpenStudy (dls):

x=Asin(2 pi/T)t for x=A.. A=Asint(2 pi/T)t (2pi/T)t=pi/2 so..t=T/4 and.. when at x=A/2 A/2=Asin(2pi/T)t'=1/2 (2pi/T)t'=pi/6 OR 5pi/6 we discard pi/6 and proceed with 5pi/6 WHY?

OpenStudy (shubhamsrg):

how do you say we discard pi/6 mathematically, range of sin^-1 is from -pi/2 to pi/2 only

OpenStudy (dls):

idk solution says "Neglected" to pi/6 and proceeds with 5pi/6 jab ata hai to pi/6 phase hota hai return karega to 5pi/6 hoga idk something like that no idea..

OpenStudy (dls):

@Vincent-Lyon.Fr !!

hartnn (hartnn):

yeah,